OpenStudy (anonymous):

If A=2 and B=6 and you get 50 how did A and B make 50?

OpenStudy (anonymous):

50 = 2A + B6

OpenStudy (anonymous):

thats the equation but how did you get 50 show the numbers that make 50 using 2 and 6

OpenStudy (anonymous):

Well I guessed 10 as A first. That would leave: 2*(10) + 6B = 50 20 + 6B = 50 6B= 50 - 20 6B = 30 B= 5

OpenStudy (anonymous):

It's mostly just a guess and check problem

OpenStudy (anonymous):

great answer just as i expected